34+21=20
69+35=46
31+29=?

Answer by hoomanc(14)

It really drives me crazy for last hours,
(sumation of digits x 2)
34+21=20 >>> (3+4+2+1)*2
69+35=46 >>> (6+9+3+5)*2
31+29=? >>> (3+1+2+9)*2 >>> 30
but seriously i was totally disappointed by myself!!!!